OpenStudy (anonymous):

Find exact value of the expression find sin (2theta) , when cos(theta) =4/5 and theta lies in quadrant IV

OpenStudy (anonymous):

To start with the triangle is a 3,4,5 triangle based on cos theta = 4/5. Secondly the sin (2theta) is equal to 2sin(theta)cos(theta)[double angle identity]. The sin of theta would equal 3/5. So 2(3/5)(4/5) = 24/5.

OpenStudy (anonymous):

thanx for your help